# Clear Minds Think Alike: What Makes LLM Fine-tuning Robust? A Study of Token Perplexity

### v1

@@ 8 Conclusion
Appendices
Appendix A Other LLM series performance
Similar to Table 2, we also have performances of other models. Due to space limitation, we add the results of Mistral-7B Instruct model of Self-Output, Rephrase, Ground Truth, and STM performances here as Table 8 shows: our conclusion that Self-Output and STM strategies for training data generation still holds for different model sizes and series.
Table 8: Performance for Ground Truth, Rephrasal, Self-Output, and STM methods across datasets MBPP and MATH. Values of a cell block in gray represent the in-domain testing performance of the model trained with in-domain task data.
Model in-domain method MBPP MATH ARC gsm8k BIRD
task
Gemma2-2B - base model 0.5106 0.2810 0.7474 0.5845 0.1389
@@ Appendix A Other LLM series performance
stm 0.4921 $\downarrow$ 0.2980 $\uparrow$ 0.7483 $\uparrow$ 0.5641 $\downarrow$ 0.1375 $\downarrow$
@@
Mistral-7B - base model 0.4868 0.1720 0.6408 0.3169 0.1284
@@
Llama3-8B - base model 0.5926 0.3140 0.7816 0.7255 0.2053
@@
A.1 STM performance on different models
@@
Table 9: STM performance comparison between Gemma2-2B Instruct, Mistral-7B-v0.3 Instruct and Llama3-8B Instruct. The percentage indicates the number of filtered tokens from the training data.
@@
A.2 Optimal threshold selection of Gemma2-2B for STM
In addition to Self-Output results, we also examine the effectiveness of STM selection strategies with Gemma Model as Figure 5 shows. Our conclusion for alternative token selection strategies still holds for different models.
Figure 5: MATH SFT using STM method on Gemma2-2B Instruct on different levels of token filtering levels. The best in domain performance also matches peak performance on out of domain tasks : GSM8k, ARC, MBPP as well.
A.3 alternative STM strategies of Llama3-8B Instruct.
We also conduct the STM strategires on Llama3-8B Instruct model (our best model series so far). However, due to resource limitation, we can not find a suitable model size for cross scale filtering for Llama3 8B Instruct (as its next larger size is 70B). As Figure 10 shows, we still have pure STM strategy as the best masking method for the best in-domain and out-domain task performance.
Table 10: List of alternative token selection methodology supervised fine-tuned on MATH in Llama3-8B Instruct. Here we only show the STM, DPF and STM+DPF version since Llama3-8B does not have a feasible model size to implement the cross-scale setting for comparison.
Llama3-8B Filter MATH MBPP GSM8K ARC BIRD
@@ A.2 Optimal threshold selection of Gemma2-2B for STM
A.4 convergence curve with Gemma2-2B Instruct
To validate the model size effect in terms of model convergence, we also observe the training procedure of Gemma2-2B Instruct. As Figure 6 shows, Self-Output and STM settings have earlier convergence time step in terms of performance and validation loss. And maintain a relative low training loss curve compared to Ground Truth data training as the same trend we see in Llama3 8B’s case.
Figure 6: MBPP in-domain testing accuracy, validation and training loss of ground truth, Self-Output, and STM data for Gemma2-2B Instruct.
Appendix B High perplexity tokens filtered by STM
@@
Appendix C Reproducibility
C.1 Training data cost
As we know there is efficiency difference of the data preparation between Self-Output, Rephrase and Selective Token Masking/Ground Truth. To brief about the difference. Self-Output is the most resource-exhausted since it requires to generate $N$ samples (usually set 32) and validate the correctness of samples (including parsing out the answer, and match with gold label or pass test cases), then do forward pass on those samples to pick the lowest perplexity data. On the other hand, STM only require a single forward pass on Ground Truth data to calculate perplexity of each token to do the masking in training peoxwaa. As Table LABEL:tab:efficiency shows, for a simple MBPP task, on a A100 GPU server to prepare for 374 instances takes about more than 18 gpu hours to complete the process of Self-Output to generate trainable data. While STM only requires 4 minutes to calculate token perplexity.
Table 11: Comparison of efficiency between Self-Output and STM in terms of gpu hours
Model Name Task Name Number of Sample Overall Correctness GPU hours
Ground Truth MBPP 374 100% 0
STM Llama3-8B Insruct MBPP 374 100% 4.5 minutes
Self-Output Llama3-8B Instruct MBPP 213 57% 16.8 hours
Rephrase Llama3-8B Instruct MBPP 327 87.4% 36.8 minutes
C.2 Model Training Resource
We train meta-llama/Meta-Llama-3-8B-Instruct, unsloth/mistral-7b-instruct-v0.3 and google/gemma-2-2b-it using two NVIDIA A100 Tensor Core GPU with VRAM 40GB for each GPU and additional RAM of 96GB. In terms of gpu hour, Our training experiments takes at most 2 hours for MBPP task, 4 to 6 hours for MATH task. As for the evaluation experiments for other out-domain tasks, evaluation in ARC-Challenge takes 3.5 hours, evaluation in GSM8k takes 4.5 hours, and evaluation in BIRD task takes 7.5 hours.
C.3 Evaluation Prompt
@@
In MBPP task, we combine a problem description (‘text‘), an answer format instruction and testing cases (‘test_list‘) into the user prompt. The generated codes will be executed with the test case to match the execution results of ground truth codes. The following TextBox demonstrate an example of MBPP data prompts:
@@
In MATH task, we add a prefix instruction before the problem description (‘problem‘) as the inference prompt. The answer will be parsed from the generated output using format of ‘\$ANSWER‘. The following TextBox demonstrate and example of MATH data prompts:
MATH prompt Solve the following math problem step by step. The last line of your response should be of the form Answer: \$ANSWER (without quotes) where \$ANSWER is the answer to the problem. problem Remember to put your answer on its own line after "Answer:", and you do not need to use a boxed command.
For OOD task,
In ARC-Challenge task, we add a Question-Answering task instruction and an answer format instruction before the ‘question‘ and answer choices (‘choices‘) as the user role prompt, and directly use the ‘answerKey‘ value to match with the parsed results from the output. The TextBox shows the user prompt is like:
@@
In GSM8k task, we add a Question-Answering task instruction and an answer format instruction before the ’question’ as the user role prompt. To evaluate the output, we directly parse the format of ’####$\#\#\#\#<number\,\,only\,\,answer>$’ from the generated results, and math it with the ’answer’ value from the dataset. The TextBox shows the user prompt is like:
@@
Appendix D Other STM alternatives
Besides cross-scale setting, inspired by [15, 17], we also investigate possible STM settings that acquire a model trained on ground truth data: Differential Perplexity Filtering (DPF): This approach leverages the perplexity changes induced by model training on unfiltered ground truth data. By computing the perplexity differential between the base and fine-tuned models, we identify tokens that demonstrate improved learnability during unrestricted training. The final training set is then constructed with tokens that exhibited reduced perplexity scores after fine-tuning, which potentially captures naturally learnable patterns in the data.
STM with DPF: This method extends the Differential Perplexity Filtering by incorporating our STM threshold-based masking mechanism. By combining both approaches, we aim to benefit from both the identification of learnable tokens through SFT and the prevention of high-perplexity token influence through STM, potentially offering a more robust token selection strategy than either method alone.
Table 12 shows the results for the best-performance models of the original STM and the other alternative of STM. We found that using DPF directly or applying STM on DPF would mask a much higher rate of tokens in training data, which leads to a worse training quality in terms of in-domain and out-of-domain tasks performances. One setting different from the previous works is that we did not prepare a high quality or high scored data for the reference model to train. Instead, we use ground truth data to filter the "unlearnable" tokens out. However, even though we mask out unlearnable tokens out, the remaining unfiltered tokens are still too few to maintain good quality of training.
Table 12: List of alternative token selection methodology supervised fine-tuned on MATH in Gemma2-2B Instruct. Here we choose Gemma2 series because Gemma2 has feasible model sizes to implement the cross-scale setting for comparison.
Gemma2-2B Filter MATH MBPP GSM8K ARC BIRD
@@ Appendix D Other STM alternatives
Appendix E Correctness experiment with other models
@@
Appendix F Regularization effects on OOD
@@
We apply a range for searching best combination of weight decay and dropout rate hyperparameters when training an MBPP task on Llama3-8B-Instruct model. The weight decay is set from 0.0 to 0.5 with interval 0.1, and dropout rate is set from 0.05 to 0.5 with an interval of 0.05 to 0.1. And we pick the top 3 best performing weight decay value with its dropout rate, and choose the best 2 performances among the checkpoints of 50 steps to 300 steps (about 5 to 6 epochs training) as our results. Our experiment results as Figure 10 show that, when applying dropout or weight decay on ground truth data, the model trained with in-domain task like MBPP still overfits on training data and the testing performance degrades. On the other hand, the out-domain performance of MATH as Figure 11 shows, we found that some of the results perform better than the original Llama in out-domain task, but their in-domain task performances are found worse than the base model’s performance. It implies that the widely adopted regularization techniques are hard to optimized for both in-domain task and out-domain task performance.
@@
Figure 10: The in-domain performance of all combinations of regularization parameters on MBPP ground-truth data model (based on Llama3-8B-Instruct) and comparison between the original Llama3-8B-Instruct model’s performance on MBPP testing dataset.
@@
Figure 11: The out-domain MATH performance of all combinations of regularization parameters on MBPP ground-truth data model (based on Llama3-8B-Instruct) and comparison between the original Llama3-8B-Instruct model’s performance on MATH testing dataset.
Appendix G Rank analysis on LoRA weights
@@
To further understand this behavior, we examine the relationship between weight updates and the information-theoretic properties of the training data. For a given sequence of tokens, we define the surprisal as $I(x,y)=-\sum_{t=1}^{T}\log P(y_{t}|x_{<t};\theta)$, where $P(y_{t}|x_{<t};\theta)$ represents the model’s predicted probability for the true token $y_{t}$ given the preceding context. In LoRA, weight updates for each transformer layer are parameterized as low-rank matrices $\Delta W=BA$, where $B\in\mathbb{R}^{d\times r}$, $A\in\mathbb{R}^{r\times k}$, and $r\ll\min(d,k)$. Through Singular Value Decomposition (SVD), $\Delta W=U\Sigma V^{\top}$, we can analyze both the magnitude of singular values in $\Sigma$ and its effective rank given a threshold $\tau$.
@@
We hypothesize that training on low-surprisal data (Self-Output or Rephrase) requires smaller weight adjustments compared to high-surprisal data (Ground Truth). This is because high-surprisal sequences induce gradients that differ significantly from those generated by data the model is already familiar with, necessitating adjustments along more independent directions in the parameter space. Our empirical analysis, comparing LoRA weight updates across training conditions with identical hyperparameters and training steps, confirms this hypothesis: Ground Truth data consistently produces updates with higher effective rank, as evidenced by the singular value spectrum of $\Delta W$. This aligns with our theoretical framework, where high-surprisal data requires the model to adjust its parameters along more dimensions to minimize the loss for these unexpected sequences.
@@
Figure 12: Effective singular value rank on layers 17th to 32th on self-attention output projection LLaMA 3 8B LoRA weights finetuned on MBPP at step 164
@@
Figure 13: Singular values on layers 17th to 32th on self-attention output projection LoRA weights finetuned on MBPP at step 164

### v2

@@ Appendix B Other LLM series performance
OLMo 2 7B Instruct
Baseline Fine-tuning -11.64 -9.05
$STM_{\tau=2.5}$ (25.83%) -6.12 -0.50
Gemma 2 IT 9B
Baseline Fine-tuning 7.14 4.67
$STM_{\tau=2.5}$ (20.84%) 13.49 2.58
The results of OLMo 2 7B Instruct imply that with such STM setting, the non-target task performance is still comparable to the original model, and the Gemma 2 IT 9B results indicate that with comparable improvement on non-target tasks after finetuning using either baseline finetuning or STM, STM brings more significant improvement on the target task. Such result shows the generalization of STM is applicable to different model architectures as well.
Table 10: Performance for Ground Truth, Rephrase, Self-Output, and STM methods across datasets MBPP and MATH. Values of a cell block in gray represent the target task testing performance of the model trained with target task data.
Model target task method MBPP MATH ARC gsm8k BIRD
Gemma 2 IT 2B - base model 0.5106 0.2810 0.7474 0.5845 0.1389
@@
stm 0.5106 0.2980 $\uparrow$ 0.7483 $\uparrow$ 0.5641 $\downarrow$ 0.1375 $\downarrow$
@@
Mistral 7B-Instruct - base model 0.4868 0.1720 0.6408 0.3169 0.1284
@@
Llama 3 8B Instruct - base model 0.5926 0.3140 0.7816 0.7255 0.2053
@@
B.1 STM performance on different models
@@
Table 11: STM performance comparison between Gemma 2 IT 2B Instruct, Mistral 7B-Instruct and Llama3 8B Instruct fine-tuned on MATH target task. The percentage indicates the number of filtered tokens from the training data.
@@
B.2 Optimal threshold selection for STM
In addition to Self-Output results, we also examine the effectiveness of STM selection strategies with Llama and Gemma Model as Figure 5 and Figure 6 shows. Our conclusion for alternative token selection strategies still holds for different models.
Figure 5: MATH SFT using STM method on Llama 3 8B Instruct on different levels of token filtering levels. The best in domain performance also matches peak performance on out of domain tasks : GSM8k, ARC, MBPP as well.
Figure 6: MATH SFT using STM method on Gemma 2 IT 2B on different levels of token filtering levels. The best in domain performance also matches peak performance on out of domain tasks : GSM8k, ARC, MBPP as well.
B.3 Alternative STM strategies of Llama 3 8B Instruct.
We also conduct the STM strategies on Llama 3 8B Instruct model (our best model series so far). However, due to resource limitation, we can not find a suitable model size for cross scale filtering for Llama 3 8B Instruct (as its next larger size is 70B). As Table 12 shows, we still have pure STM strategy as the best masking method for the best target and non-target task performance.
Table 12: List of alternative token selection methodology supervised fine-tuned on MATH in Llama 3 8B Instruct. Here we only show the STM, DPF and STM+DPF version since Llama 3 8B Instruct does not have a feasible model size to implement the cross-scale setting for comparison.
Llama 3 8B Instruct Filter MATH MBPP GSM8K ARC BIRD
@@ B.2 Optimal threshold selection for STM
B.4 Other fine-tuning techniques applied with STM on Llama 3 8B Instruct
We have done experiments of STM applying with LLM adaptation like LoRA and DoRA on Llama 3 8B Instruct as Table 13 shows. Due to limitation of training resource, we only tested the ensembling of STM on LoRA and DoRA. On both adapters, STM enhances the TI and BWT performances.
Table 13: STM applied on DoRA and LoRA with MBPP on Llama 3 8B Instruct.
[…]
LoRA -2.23 -34.71
LoRA + $STM_{\tau=2.5}$ 3.2 -0.3
DoRA -8.03 -1.89
DoRA + $STM_{\tau=2.5}$ 3.58 1.69
B.5 Convergence curve with Gemma 2 IT 2B
To validate the model size effect in terms of model convergence, we also observe the training procedure of Gemma 2 IT 2B. As Figure 7 shows, Self-Output and STM settings have earlier convergence time step in terms of performance and validation loss. And maintain a relative low training loss curve compared to Ground Truth data training as the same trend we see in Llama 3 8B Instruct’s case.
Figure 7: MBPP target task testing accuracy, validation and training loss of ground truth, Self-Output, and STM data for Gemma 2 IT 2B.
Appendix C High perplexity tokens filtered by STM
@@
Appendix D Reproducibility
D.1 Training data cost
As we know there is efficiency difference of the data preparation between Self-Output, Rephrase and Selective Token Masking/Ground Truth. To brief about the difference, Self-Output is the most resource-exhausted since it requires to generate $N$ samples (usually set 32), validate the correctness of samples (including parsing out the answer, and match with gold label or pass test cases), and then do forward passes on those samples to pick the lowest perplexity sentence as training data. On the other hand, STM only requires a single forward pass on Ground Truth data to calculate perplexity of each token to do the masking in training process. As Table 14 shows, for a simple MBPP task, on a A100 GPU server to prepare for 374 instances (train) + 90 instances (validation) for $N$=32 takes about more than 18 gpu hours to complete the process of Self-Output to generate trainable data. While STM only requires 4 minutes to calculate the whole token perplexity.
Table 14: Comparison of efficiency between STM, Self-Output and Rephrase in terms of gpu hours on MBPP target task.
Model Name Number of Sample Correctness GPU hours
Ground Truth 374 100% 0
STM Llama3-8B Insruct 374 100% 4.5 minutes
Self-Output Llama3-8B Instruct 213 57% 16.8 hours
Rephrase Llama3-8B Instruct 327 87.4% 36.8 minutes
D.2 Model training resource
We train Llama 3 8B Instruct, Mistral 7B-Instruct and Gemma 2 IT 2B using two NVIDIA A100 Tensor Core GPU with VRAM 40GB for each GPU and additional RAM of 96GB. In terms of gpu hour, Our training experiments takes at most 2 hours for MBPP task, 4 to 6 hours for MATH task. As for the evaluation experiments for other non-target tasks, evaluation in ARC-Challenge takes 3.5 hours, evaluation in GSM8k takes 4.5 hours, and evaluation in BIRD task takes 7.5 hours.
D.3 Evaluation prompt
@@
In MBPP task, we combine a problem description (‘text‘), an answer format instruction and testing cases (‘test_list‘) into the user prompt. The generated codes will be executed with the test case to match the execution results of ground truth codes. The following TextBox demonstrates an example of MBPP data prompts:
@@
In MATH task, we add a prefix instruction before the problem description (‘problem‘) as the inference prompt. The answer will be parsed from the generated output using format of ‘\$ANSWER‘. The following TextBox demonstrates an example of MATH data prompts:
MATH prompt user prompt Solve the following math problem step by step. The last line of your response should be of the form Answer: \$ANSWER (without quotes) where \$ANSWER is the answer to the problem. problem Remember to put your answer on its own line after "Answer:", and you do not need to use a boxed command.
For Non-target task: In ARC-Challenge task, we add a Question-Answering task instruction and an answer format instruction before the ‘question‘ and answer choices (‘choices‘) as the user role prompt, and directly use the ‘answerKey‘ value to match with the parsed results from the output. The TextBox shows the user prompt is like:
@@
In GSM8k task, we add a Question-Answering task instruction and an answer format instruction before the ’question’ as the user role prompt. To evaluate the output, we directly parse the format of ’####$\#\#\#\#<number\,\,only\,\,answer>$’ from the generated results, and match it with the ’answer’ value from the dataset. The TextBox shows an example of the user prompt:
@@
Appendix E Other STM alternatives
Besides cross-scale setting, inspired by [19, 21], we also investigate possible STM settings that acquire a model trained on ground truth data:
Differential Perplexity Filtering (DPF). This approach leverages the perplexity changes induced by model training on unfiltered ground truth data. By computing the perplexity differential between the base and fine-tuned models, we identify tokens that demonstrate improved learnability during unrestricted training. The final training set is then constructed with tokens that exhibited reduced perplexity scores after fine-tuning, which potentially captures naturally learnable patterns in the data.
STM with DPF. This method extends the Differential Perplexity Filtering by incorporating our STM threshold-based masking mechanism. By combining both approaches, we aim to benefit from both the identification of learnable tokens through SFT and the prevention of high-perplexity token influence through STM, potentially offering a more robust token selection strategy than either method alone.
Table 15 shows the results for the best-performance models of the original STM and the other alternative of STM. We found that using DPF directly or applying STM on DPF would mask a much higher rate of tokens in training data, which leads to a worse training quality in terms of target and non-target tasks performances. One setting different from the previous works is that we did not prepare a high quality or high scored data for the reference model to train. Instead, we use ground truth data to filter the "unlearnable" tokens out. However, even though we mask out unlearnable tokens out, the remaining unfiltered tokens are still too few to maintain good quality of training.
Table 15: List of alternative token selection methodology supervised fine-tuned on MATH in Gemma 2 IT 2B . Here we choose Gemma 2 series because Gemma 2 has feasible model sizes to implement the cross-scale setting for comparison.
Gemma 2 IT 2B Filter MATH MBPP GSM8K ARC BIRD
@@ Appendix E Other STM alternatives
${DPF}$ 39.2% 22.5 40.2 1.1 6.0 11.0
${\tau=0}$
${STM+DPF}$ 44.7% 17.3 43.7 9.1 14.4 11.7
${\tau=100}$
Appendix F Correctness experiment with other models
@@
Appendix G Regularization effects on non-target tasks
@@
We apply a range for searching best combination of weight decay and dropout rate hyperparameters when training an MBPP task on Llama3-8B-Instruct model. The weight decay is set from 0.0 to 0.5 with interval 0.1, and dropout rate is set from 0.05 to 0.5 with an interval of 0.05 to 0.1. And we pick the top 3 best performing weight decay value with its dropout rate, and choose the best 2 performances among the checkpoints of 50 steps to 300 steps (about 5 to 6 epochs training) as our results. Our experiment results as Figure 11 show that, when applying dropout or weight decay on ground truth data, the model trained with target task like MBPP still overfits on training data and the testing performance degrades. On the other hand, the non-target performance of MATH as Figure 12 shows, we found that some of the results perform better than the original Llama in non-target task, but their target task performances are found worse than the base model’s performance. It implies that the widely adopted regularization techniques are hard to optimized for both target task and non-target task performance.
@@
Figure 11: The target task performance of all combinations of regularization parameters on MBPP ground-truth data model (based on Llama3-8B-Instruct) and comparison between the original Llama3-8B-Instruct model’s performance on MBPP testing dataset.
@@
Figure 12: The non-target MATH task performance of all combinations of regularization parameters on MBPP ground-truth data model (based on Llama3-8B-Instruct) and comparison between the original Llama3-8B-Instruct model’s performance on MATH testing dataset.
Appendix H Rank analysis on LoRA weights
@@
To further understand this behavior, we examine the relationship between weight updates and the information-theoretic properties of the training data. For a given sequence of tokens, we define the token perplexity as $I(x,y)=-\sum_{t=1}^{T}\log P(y_{t}|x_{<t};\theta)$, where $P(y_{t}|x_{<t};\theta)$ represents the model’s predicted probability for the true token $y_{t}$ given the preceding context. In LoRA, weight updates for each transformer layer are parameterized as low-rank matrices $\Delta W=BA$, where $B\in\mathbb{R}^{d\times r}$, $A\in\mathbb{R}^{r\times k}$, and $r\ll\min(d,k)$. Through Singular Value Decomposition (SVD), $\Delta W=U\Sigma V^{\top}$, we can analyze both the magnitude of singular values in $\Sigma$ and its effective rank given a threshold $\tau$.
@@
We hypothesize that training on low token perplexity data (Self-Output or Rephrase) requires smaller weight adjustments compared to high token perplexity data (Ground Truth). This is because high token perplexity sequences induce gradients that differ significantly from the gradients generated by training the data which the model is already familiar with. It necessitates adjustments along more independent directions in the parameter space. Our empirical analysis, comparing LoRA weight updates across training conditions with identical hyperparameters and training steps, confirms this hypothesis: Ground Truth data consistently produces updates with higher effective rank, as evidenced by the singular value spectrum of $\Delta W$. This aligns with our theoretical framework, where high perplexity tokens require the model to adjust parameters along more dimensions to minimize the loss for these unexpected sequences. In Figure 13 and Figure 14, we identify effective singular value ranks on layers 17th to 32nd on self-attention output projection of Llama 3 8B Instruct LoRA weights and find that Ground Truth fine-tuning enables more ranks to learn with higher values for each attention, indicating more perturbations on training to degrade the general capabilities to fit a target task.
@@
Figure 13: Effective singular value rank on layers 17th to 32nd on self-attention output projection Llama 3 8B Instruct LoRA weights finetuned on MBPP at step 164
@@
Figure 14: Singular values on layers 17th to 32tnd on self-attention output projection LoRA weights finetuned on MBPP at step 164